OpenStudy (anonymous):

someone check my answer plz If the coordinates of the points D, E, F, G, and H are -3, -6, 8, 11, and 13, respectively, which segment is the longest? DE EG FG FH<--- my answer

jimthompson5910 (jim_thompson5910):

what did you get for DE, EG, FG and FH?

jimthompson5910 (jim_thompson5910):

you subtract the values to find the lengths of the segments

jimthompson5910 (jim_thompson5910):

example: from F to G = 8-11 = -3 take the absolute value: |-3| = 3 FG is 3 units long

OpenStudy (anonymous):

DE <--- -3- -6=3 EG <-- -6 - 8= -14 FG<-- 8 - 11=-3 FH<---- 8-13=-5

jimthompson5910 (jim_thompson5910):

EG should be from -6 to +11

jimthompson5910 (jim_thompson5910):

so either -6 - 11 = -17 |-17| = 17 or 11 - (-6) = 11+6 = 17

jimthompson5910 (jim_thompson5910):

EG is 17 units long

OpenStudy (anonymous):

is then is will eg is the longest

OpenStudy (anonymous):

eg is the answer cuz its the longest

jimthompson5910 (jim_thompson5910):

correct